[Brahms-l] Shift report 20040205 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Thu Feb 05 2004 - 23:55:43 EST
------------------------------------------------------------
Date: Thu Feb  5 23:55:42 2004, Shift: 16:00-24:00
Supervisor: EBJ, with: 
Log book#15 pages 49 - 49

Shift Summary:
------------------------------------------------------------
1.  RHIC has only begun to store beam at the end of this shift.  They had probelm all night.

2. I only took the pedestal runs.

3. The checklist indicates that the detector has been running fine.  I just ramped up the magnets, and everthing went well with that.  I haven't noticed any computer problems either.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu Feb 5 23:56:11 2004

This archive was generated by hypermail 2.1.8 : Thu Feb 05 2004 - 23:56:30 EST